Compute the value of the discriminant and give the number or real solutions to the quadratic equation.
-9x^2 +6x -1 =0

Answer by mukhopadhyay(490)

Discriminant for quadratic equation ax^2+bx+c =0 is (b^2-4ac)
Here a=-9, b=6 and c=-1
(b^2-4ac) = 36-4(-9)(-1) = 36-36 = 0
Since the discriminant is zero, there is only one real solution to this equation, which is -b/2a = 6/18 = 1/3
Answer: x= 1/3